OpenStudy (pleasehelpme):

please solve this problem X^2-6x+8=0

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=X^2-6x%2B8 The integer roots are: x = 2 (3 n^2-2 n+1), X = -2 (3 n-1), n element Z